# DeFoG: Discrete Flow Matching for Graph Generation

### v1

@@ E.3 Hyperparameter Tuning
Table 6: Training and sampling parameters for full-step sampling (500 or 1000 steps for synthetic and molecular datasets respectively).
@@
Dataset Initial Distribution Train Distortion Sample Distortion $\omega$ (Target Guidance) $\eta$ (Stochasticity) Exact Exp $\gamma$
Planar Marginal Identity Polydec 0.05 50 True
Tree Marginal Polydec Polydec 0.0 0.0 False
SBM Absorbing Identity Identity 0.0 0.0 False
QM9 Marginal Identity Polydec 0.0 0.0 False
QM9(H) Marginal Identity Polydec 0.05 0.0 False
Moses Marginal Polydec Polydec 0.5 200 False
Guacamol Marginal Polydec Polydec 0.1 300 False
TLS Marginal Identity Polydec 0.05 0.0 False 2.0
@@
Appendix F Additional Results
In this section, we present additional experimental results. We begin with the complete tables showcasing results for synthetic datasets in Sec. F.1. Next, Sec. F.2 focuses on molecular generation tasks, including results for the QM9 dataset and the full tables for MOSES and Guacamol. Finally, in Sec. F.3, we analyze the time complexity of various additional features that enhance expressivity in graph diffusion models.
F.1 Synthetic Graph Generation
@@
Table 7: Graph Generation Performance on Synthetic Graphs. We present the results of DeFoG across five sampling runs, each generating 40 graphs, reported as mean ± standard deviation. The remaining values are obtained from Bergmeister et al. (2023). Additionally, we include results for Cometh (Siraudin et al., 2024) and DisCo (Xu et al., 2024). For the average ratio computation, we adhere to the method outlined by Bergmeister et al. (2023), excluding statistics where the training set MMD is zero.
@@
F.2 Molecular Graph Generation
@@
Table 8: Molecule generation on QM9. We present the results over five sampling runs of 10000 generated graphs each, in the format mean ± standard deviation.
@@
Model Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
Training set 99.3 99.2 0.03 97.8 99.9 0.01
@@
GraphNVP (Madhawa et al., 2019) $83.1$ 99.2
GDSS (Jo et al., 2022) $95.7$ $\underline{98.5}$ $2.9$
DiGress (Vignac et al., 2022) 99.0±0.0 96.2±0.2 95.4±1.1 97.6±0.4
@@
DeFoG (# sampling steps = 50) 98.9±0.1 96.2±0.2 0.26±0.00 97.1±0.0 94.8±0.0 0.31±0.00
DeFoG (# sampling steps = 500) 99.3±0.0 96.3±0.3 0.12±0.00 98.0±0.0 96.7±0.0 0.05±0.00
Additionally, we provide the complete version of Tab. 2, presenting the results for MOSES and Guacamol separately in Tab. 9 and Tab. 10, respectively. We include models from classes beyond diffusion models to better contextualize the performance achieved by DeFoG.
@@
LSTM (Graves & Graves, 2012) Smiles 95.9 95.9 87.4 99.1 91.3
@@
F.3 Efficiency of Additional Features
In graph diffusion methods, the task of graph generation is decomposed into a mapping of a graph to a set of marginal probabilities for each node and edge. This problem is typically addressed using a Graph Transformer architecture, which is augmented with additional features to capture structural aspects that the base architecture might struggle to model effectively (Vignac et al., 2022; Xu et al., 2024; Siraudin et al., 2024) otherwise. In this section, we conduct a time complexity analysis of the additional features commonly employed. Specifically, we compare the spectral encodings and cycle encodings (up to $6$-cycles) proposed by Vignac et al. (2022) to the RRWP encodings (12 steps) from Siraudin et al. (2024) (originally from Ma et al. (2023)). While both cycle and RRWP encodings primarily involve matrix multiplications, spectral encodings necessitate more complex algorithms for eigenvalue and eigenvector computation. As shown in Tab. 11, cycle and RRWP encodings are more computationally efficient, particularly for larger graphs where eigenvalue computation becomes increasingly costly. These results support the use of RRWP encodings over the combined utilization of cycle and spectral features.
For the graph sizes considered in this work, the additional feature computation time remains relatively small compared to the model’s forward pass and backpropagation. However, as graph sizes increase — a direction beyond the scope of this paper — this computational gap could become significant, making RRWP a proper encoding for scalable graph generative models.
Table 11: Computation time for different additional features.

### v3

@@ F.4 Hyperparameter Tuning
Table 5: Training and sampling parameters for full-step sampling (500 or 1000 steps for molecular and synthetic datasets, respectively).
@@
Dataset Initial Distribution Train Distortion Sample Distortion $\omega$ (Target Guidance) $\eta$ (Stochasticity) $\gamma$
Planar Marginal Identity Polydec 0.05 50
Tree Marginal Polydec Polydec 0.0 0.0
SBM Absorbing Identity Identity 0.0 0.0
QM9 Marginal Identity Polydec 0.0 0.0
QM9(H) Marginal Identity Polydec 0.05 0.0
Moses Marginal Polydec Polydec 0.5 200
Guacamol Marginal Polydec Polydec 0.1 300
TLS Marginal Identity Polydec 0.05 0.0 2.0
ZINC250k Marginal Identity Polydec 0.0 200
@@
Appendix G Additional Results
First, we discuss conditional generation for real-world digital pathology graph generation. Next, we provide comprehensive tables detailing results on synthetic datasets in Sec. G.2. Then, Sec. G.3 focuses on molecular generation tasks, including results for the QM9 dataset and the full tables for MOSES and Guacamol. Finally, in Sec. G.4, we analyze the time complexity of various additional features that enhance expressivity in graph diffusion models.
G.1 Conditional Generation
Setup
Tertiary Lymphoid Structure (TLS) graph datasets have been recently released with graphs built from digital pathology data (Madeira et al., 2024). These graphs are split into two subsets — low TLS and high TLS — based on their TLS content, a biologically informed metric reflecting the cell organization in the graph structure. Previously, models were trained and evaluated separately for each subset. To demonstrate the flexibility of DeFoG, we conditionally train it across both datasets simultaneously, using the low/high TLS content as a binary label for each graph. More details on the used conditional framework in Appendix E.
Table 6: TLS conditional generation results.
TLS Dataset
Model V.U.N. $\uparrow$ TLS Val. $\uparrow$
Train set 0.0 100
GraphGen (Goyal et al., 2020) 40.2±3.8 25.1±1.2
BiGG (Dai et al., 2020) 0.6±0.4 16.7±1.6
SPECTRE (Martinkus et al., 2022) 7.9±1.3 25.3±0.8
DiGress+ (Madeira et al., 2024) 13.2±3.4 12.6±3.0
ConStruct (Madeira et al., 2024) 99.1±1.1 92.1±1.3
DeFoG (# steps = 50) 44.5±4.2 93.0±5.6
DeFoG (# steps = 1,000) 94.5±1.8 95.8±1.5
Results
G.2 Synthetic Graph Generation
@@
Table 7: Graph Generation Performance on Synthetic Graphs. We present the results of DeFoG across five sampling runs, each generating 40 graphs, reported as mean ± standard deviation. The remaining values are obtained from Bergmeister et al. (2023). Additionally, we include results for Cometh (Siraudin et al., 2024) and DisCo (Xu et al., 2024). For the average ratio computation, we adhere to the method outlined by (Bergmeister et al., 2023), excluding statistics where the training set MMD is zero.
@@
G.3 Molecular Graph Generation
@@
Table 8: Molecule generation on QM9. We present the results over five sampling runs of 10000 generated graphs each, in the format mean ± standard deviation. We include the results of Relaxed Validity, which accounts for charged molecules to facilitate comparison, as different methods may report varying types of validity.
@@
Model Valid $\uparrow$ Relaxed Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Relaxed Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
Training set 99.3 99.5 99.2 0.03 97.8 98.9 99.9 0.01
@@
GraphNVP (Madhawa et al., 2019) $83.1$ $\underline{99.2}$
GDSS (Jo et al., 2022) $95.7$ $98.5$ $2.9$
DiGress (Vignac et al., 2022) 99.0±0.0 96.2±0.1 95.4±1.1 97.6±0.4
GruM(Jo et al., 2024) 99.2 96.7 0.11
CatFlow(Eijkelboom et al., 2024) 99.8 100.0 0.44
@@
DeFoG (# sampling steps = 50) 98.9±0.1 99.2±0.0 96.2±0.2 0.26±0.00 97.1±0.0 98.1±0.0 94.8±0.0 0.31±0.00
DeFoG (# sampling steps = 500) 99.3±0.0 99.4±0.1 96.3±0.3 0.12±0.00 98.0±0.0 98.8±0.0 96.7±0.0 0.05±0.00
Additionally, we provide the complete version of Tab. 2, presenting the results for MOSES, Guacamol and ZINC250k datasets separately in Tab. 9, Tab. 10 and Tab. 11, respectively. We include models from classes beyond diffusion models to better contextualize the performance achieved by DeFoG. We analyze the performance of diffusion and flow-based methods on MOSES and Guacamol in the main paper (see Sec. 5.1). Here, we focus on the same analysis for ZINC250k. As shown in Tab. 11, DeFoG achieves state-of-the-art performance on this benchmark. Notably, it also attains superior FCD scores with only 50 sampling steps, surpassing existing diffusion and flow-based methods.
@@
LSTM (Hochreiter & Schmidhuber, 1997) Smiles 95.9 95.9 87.4 99.1 91.3
@@
Table 11: Molecular generation on ZINC250k dataset.
Model Val. $\uparrow$ Uniqueness $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Scaffold $\uparrow$
GruM (Jo et al., 2024) 98.65 2.257 0.0015 0.5299
GBD (Liu et al., 2024) 97.87 2.248 0.0018 0.5042
CatFlow (Eijkelboom et al., 2024) 99.21 100.00 13.211
DeFoG (50 steps) 96.65±0.16 99.99±0.01 2.123±0.029 0.0022±0.0001 0.4245±0.0109
DeFoG 99.22±0.08 99.99±0.01 1.425±0.022 0.0008±0.0001 0.5903±0.0099
G.4 Impact of Additional Features
In graph diffusion methods, the task of graph generation is decomposed into a mapping of a graph to a set of marginal probabilities for each node and edge. This problem is typically addressed using a Graph Transformer architecture, which is augmented with additional features to capture structural aspects that the base architecture might struggle to model effectively (Vignac et al., 2022; Xu et al., 2024; Siraudin et al., 2024) otherwise.
In this section, we evaluate the impact of using RRWP encodings as opposed to the spectral and cycle encodings (up to $6$-cycles) proposed in DiGress (Vignac et al., 2022).
In Tab. 12, we present a performance comparison of these two variants across three datasets: QM9, Planar, and SBM. The results show that RRWP achieves comparable or superior performance within the DiGress framework, validating its effectiveness as a graph encoding method. Notably, despite these improvements, DiGress’s performance remains significantly below that of DeFoG on the Planar and SBM datasets, while achieving similar validity and uniqueness on the QM9 dataset.
Table 12: Performance comparison of RRWP-based graph encoding within the DiGress framework.
Method Planar SBM QM9
V.U.N. $\uparrow$ Ratio $\downarrow$ V.U.N. $\uparrow$ Ratio $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
DiGress $77.5$ $5.1$ $60.0$ $1.7$ $99.0\pm 0.0$ $96.2\pm 0.1$
DiGress (RRWP) $90.0$ $4.0$ $70.0$ $1.7$ $99.1\pm{0.1}$ $96.6\pm{0.2}$
DeFoG (RRWP, 50 steps) $95.0\pm 3.2$ $3.2\pm 1.1$ $86.5\pm 5.3$ $2.2\pm 0.3$ $98.9\pm 0.1$ $96.2\pm 0.2$ $0.26\pm 0.00$
DeFoG (RRWP) $99.5\pm 1.0$ $1.6\pm 0.4$ $90.0\pm 5.1$ $4.9\pm 1.3$ $99.3\pm 0.0$ $96.3\pm 0.3$ $0.12\pm 0.00$
To further demonstrate the impact of using RRWP on sampling efficiency, we compare the performances of DeFoG, DiGress, and DiGress augmented with RRWP (replacing the original additional features) across a varying number of sampling steps. These results are shown in Figure 29.
(a) QM9 dataset.
(b) Planar dataset.
(c) SBM dataset.
Figure 29: Impact of RRWP features for sampling efficiency.
We observe that, while RRWP provides improvements on the Planar and SBM datasets with fewer generation steps, it is still significantly outperformed by the optimized DeFoG framework. This highlights that although RRWP is an efficient and effective graph encoding method, the primary performance gains of DeFoG stem from its continuous-time formulation featuring fully decoupled training and sampling stages.
We then perform a time complexity analysis of these methods. While both cycle and RRWP encodings primarily involve matrix multiplications, spectral encodings require more complex algorithms for eigenvalue and eigenvector computation. As shown in Tab. 13, cycle and RRWP encodings are more computationally efficient, particularly for larger graphs where eigenvalue computation becomes increasingly costly. These results also support the use of RRWP encodings over the combined utilization of cycle and spectral features.
For the graph sizes considered in this work, the additional feature computation time remains relatively small compared to the model’s forward pass and backpropagation. However, as graph sizes increase - a direction beyond the scope of this paper - this computational gap could become significant, making RRWP a suitable encoding for scalable graph generative models.
Table 13: Computation time for different additional features. The RRWP features are computed with 12 steps.